OpenStudy (anonymous):

If h(-2)=3 and h'(-2) =4 find dy/dx(h(x)/x) at x=-2?

OpenStudy (anonymous):

d/dx[h(x)/x] = (x*h' - h(x)*1)/x^2

OpenStudy (anonymous):

d/dx[h(x)/x] at x=-2 is ((-2)*(4) - (3)*1)/(-2)^2 =(-8 - 3)/4 = -11/4

OpenStudy (anonymous):

I just used the quotient rule

OpenStudy (anonymous):

can somebody check this?

OpenStudy (anonymous):

wait, i think your substitutions are wrong

OpenStudy (anonymous):

just kdiding

OpenStudy (anonymous):

@jaersyn, where?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Thank you everyone